OpenStudy (anonymous):

Please help! Simplify the complex fraction. 2/5t-3/3t / 1/2t+1/2t

OpenStudy (anonymous):

??

OpenStudy (anonymous):

What's the fraction?

OpenStudy (dan815):

hey gabby u need some help?

OpenStudy (anonymous):

yea but in english..

OpenStudy (anonymous):

I post it as a picture

OpenStudy (anonymous):

I don't see a picture.

OpenStudy (anonymous):

yea me either

OpenStudy (anonymous):

I post the fraction with the question I guess the picture didn't show up

OpenStudy (jdoe0001):

\(\large \bf \cfrac{ \frac{2}{5t}-\frac{3}{3t} }{ \frac{1}{2t}+\frac{1}{2t} }\quad ?\)

OpenStudy (anonymous):

yes!

OpenStudy (jdoe0001):

well.. got any ideas on the LCD of the numerator? \(\bf \cfrac{2}{5t}-\cfrac{3}{3t}=\cfrac{}{{\color{blue}{ lcd?}}}\)

OpenStudy (anonymous):

Would it be 2?

OpenStudy (jdoe0001):

well.... the lcd will be a expression that can be divided by 5t and also can be divided at the same time by 3t

OpenStudy (anonymous):

...15?

OpenStudy (jdoe0001):

well... yes.... sorta 15t really, notice the "t" in each term so one sec

OpenStudy (anonymous):

okay!

OpenStudy (jdoe0001):

\(\large { \bf \cfrac{ \frac{2}{5t}-\frac{3}{3t} }{ \frac{1}{2t}+\frac{1}{2t} }\implies \cfrac{ \frac{3\cdot 2-5\cdot3 }{15t} }{ \frac{1+1}{2t} }\implies \cfrac{ \frac{6-15}{15t} }{ \frac{\cancel{ 2 }}{\cancel{ 2 }t} } \\ \quad \\ \cfrac{ \frac{\cancel{ 9 }}{\cancel{ 15 }t} }{ \frac{1}{t} }\implies \cfrac{ \frac{3}{5t} }{ \frac{1}{t} }\qquad recall\implies {\color{brown}{ \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}}}\qquad thus \\ \quad \\ \cfrac{ \frac{3}{5t} }{ \frac{1}{t} }\implies \cfrac{3}{5t}\cdot \cfrac{t}{1} }\)

OpenStudy (jdoe0001):

then just multiply them and simplify, hint: you can cancel out the "t"s

OpenStudy (anonymous):

would it end up being -3/5?